# Fractionally modulated discrete Carleson’s Theorem and pointwise Ergodic Theorems along certain curves

### v1

@@ 5. Reducing the proof of Theorem 1 to the continuous case
and notice that our choice of $M$ and Lemma 4.4 from [10] give us
$\|L_{j}\|_{\ell^{1}}\lesssim 2^{-j}\sum_{2^{j-3}\leq n\leq 2^{j-1}}\min\bigg{\{}1,\frac{1}{M\|n^{c}\|}\bigg{\}}\lesssim 2^{-j}j(2^{(1-\delta_{1})j}+2^{(c/2+\delta_{1}/2)j})\lesssim 2^{-j\delta_{1}/2}+j2^{\frac{c+\delta_{1}-2}{2}j}$
and $c+\delta_{1}-2<0$ since $\delta_{1}<2-c$. Thus there exists $\rho_{1}>0$ such that $\|L_{j}\|_{\ell^{1}}\lesssim 2^{-\rho_{1}j}$. Now by defining $T_{j}f(x)\coloneqq f*L_{j}(x)$, we see that $\|T_{j}\|_{\ell^{2}\to\ell^{2}}\lesssim 2^{-\rho_{1}j}$. For $\tau\in\{+,-\}$, let
$\psi_{j}^{\tau}=\left\{\begin{array}[]{ll}\psi_{j}\bm{1}_{(0,\infty)}\text{,}&\text{if }\tau=+\\ \psi_{j}\bm{1}_{(-\infty,0)}\text{,}&\text{if }\tau=-\text{.}\\ \end{array}\right.$
@@
(5.10) $\sum_{\mathsf{\tau}\in\left\{+,-\right\}}\sup_{\lambda\in X_{j}^{c}}\sum_{|m|\leq M}\frac{1}{|m|+1}\left|\sum_{n\neq 0}f(x-n)e((\lambda\mathsf{sign}^{\mathsf{i}}(n)+m)|n|^{c})\psi_{j}^{\tau}(n)\right|\\ \lesssim\sum_{\mathsf{\tau}\in\left\{+,-\right\}}\sup_{2^{(-c+\nu)j}<|\lambda|\leq M+1/2}\left|\sum_{n\neq 0}f(x-n)e(\lambda|n|^{c})\psi_{j}^{\tau}(n)\right|\log(M+1)\\ \lesssim j\max_{\mathsf{\tau}\in\left\{+,-\right\}}\sup_{2^{(-c+\nu)j}<|\lambda|\leq 2^{\delta_{1}j}+1}\Big{|}\sum_{n\neq 0}f(x-n)e(\lambda|n|^{c})\psi_{j}^{\tau}(n)\Big{|}\text{.}$
Combining everything yields
@@
$\Big{\|}\sup_{\lambda\in X^{c}_{j}}\big{|}\big{(}\hat{f}(\cdot)m_{j}^{\mathsf{i}}(\cdot,\lambda)\big{)}^{\vee}(x)\big{|}\Big{\|}_{\ell^{2}_{dx}}\lesssim j\max_{{\mathsf{\tau}\in\left\{+,-\right\}}}\bigg{\|}\sup_{2^{(-c+\nu)j}<|\lambda|\leq 2^{\delta_{1}j}+1}\Big{|}\sum_{n\neq 0}f(x-n)e(\lambda|n|^{c})\psi_{j}^{\tau}(n)\Big{|}\bigg{\|}_{\ell^{2}_{dx}}+2^{-\rho_{1}j}\|f\|_{\ell^{2}}\text{,}$
@@
and we focus on the first term of the right hand side. We prove that for each $\tau\in\left\{+,-\right\}$ there exists $\rho_{2}>0$ such that
$\bigg{\|}\sup_{2^{\nu j}<|\lambda 2^{cj}|\leq 2^{(c+2\delta_{1})j}}\Big{|}\sum_{n\neq 0}f(x-n)e(\lambda|n|^{c})\psi_{j}^{\tau}(n)\Big{|}\bigg{\|}_{\ell^{2}_{dx}}\lesssim 2^{-\rho_{2}j}\|f\|_{\ell^{2}}\text{,}$
where for the last reduction we have assumed that $j\geq C(c)$ for some constant depending on $\delta_{1}=\delta_{1}(c)$. This is a benign assumption since we can discard any finite amount of $j$’s by (5.3). Without loss of generality we may assume that $\lambda>0$. It will be important to have further control of the scale of $2^{cj}\lambda$, so we perform a further decomposition. Note that
(5.11) $\sup_{2^{\nu j}<\lambda 2^{cj}\leq 2^{(c+2\delta_{1})j}}\Big{|}\sum_{n\neq 0}f(x-n)e(\lambda|n|^{c})\psi_{j}^{\tau}(n)\Big{|}\\ \leq\sum_{r\in\mathbb{Z}}\sup_{2^{\nu j}<\lambda 2^{cj}\leq 2^{(c+2\delta_{1})j}}\bm{1}_{[2^{r},2^{r+1})}(\lambda 2^{cj})\Big{|}\sum_{n\neq 0}f(x-n)e(\lambda|n|^{c})\psi_{j}^{\tau}(n)\Big{|}\\ \leq\sum_{\nu j-2\leq r\leq(c+2\delta_{1})j+2}\sup_{2^{r}\leq\lambda 2^{cj}<2^{r+1}}\Big{|}\sum_{n\neq 0}f(x-n)e(\lambda|n|^{c})\psi_{j}^{\tau}(n)\Big{|}\text{,}$
where for the last equality we simply note that for any $r\in\mathbb{Z}$ such that $r<\nu j-2$ or $r>(c+2\delta_{1})j+2$ we have that
$[2^{r},2^{r+1})\cap(2^{\nu j},2^{(c+2\delta_{1})j}]=\emptyset\text{.}$
@@
Note that the outer sum has $O(j)$ terms and thus to establish the desired estimate it suffices to prove that for any $r\in[\nu j-2,(c+2\delta_{1})j+2]$ we have that
$\bigg{\|}\sup_{2^{r}\leq\lambda 2^{cj}<2^{r+1}}\Big{|}\sum_{n\neq 0}f(x-n)e(\lambda|n|^{c})\psi_{j}^{\tau}(n)\Big{|}\bigg{\|}_{\ell^{2}_{dx}}\lesssim 2^{-\rho_{3}j}\|f\|_{\ell^{2}}\text{,}\quad\text{for some $\rho_{3}>0$ depending only on $c$.}$
@@
We linearize the supremum; it suffices to prove that for any measurable $\lambda_{x}\colon\mathbb{Z}\to[2^{r-cj},2^{r+1-cj})$ we have
(5.12) $\Big{\|}\sum_{n\neq 0}f(x-n)e(\lambda_{x}|n|^{c})\psi_{j}^{\tau}(n)\Big{\|}_{\ell^{2}_{dx}}\lesssim 2^{-\rho_{3}j}\|f\|_{\ell^{2}}\text{.}$
Let us fix such a function and $j\geq C(c)$. We use the $TT^{*}$ method here. It is not difficult to see that if
@@
Fix $\delta_{2}\in\big{(}0,\min\big{\{}\frac{\nu}{100},\frac{2-c}{100}\big{\}}\big{)}$. There exists $\rho_{4}>0$ such that the following estimate holds
(5.13) $|K_{\tau,j}(x,y)|\lesssim 2^{-(1+\rho_{4})j}\bm{1}_{2^{(1-\delta_{2})j}\leq|x-y|\leq 2^{j}}+2^{-j}\bm{1}_{|x-y|\leq 2^{(1-\delta_{2})j}}\text{.}$
@@
We begin with certain straightforward reductions. Firstly, we assume that $y\geq x$ and $\tau=-$ since symmetric arguments yield the result for the remaining cases. Secondly, we may assume that $|x-y|\geq 2^{(1-\delta_{2})j}$ since otherwise we clearly have
@@
Now for the case $y\geq x$ and $|x-y|\geq 2^{(1-\delta_{2})j}$ we note that
@@
[…]
@@
[…]
@@
$|f^{\prime\prime}(t)|=|c(c-1)(c-2)|\xi^{c-3}\Big{|}\lambda_{x}^{\frac{1}{c-2}}(t-x)-\lambda_{y}^{\frac{1}{c-2}}(t-y)\Big{|}$
$=|c(c-1)(c-2)|\lambda_{x}^{\frac{1}{c-2}}\xi^{c-3}\bigg{|}(t-x)-\frac{\lambda_{y}^{\frac{1}{c-2}}}{\lambda_{x}^{\frac{1}{c-2}}}(t-y)\bigg{|}\text{,}$
@@
$|f^{\prime\prime}(t)|=|c(c-1)(c-2)|\lambda_{x}^{\frac{1}{c-2}}\xi^{c-3}\big{|}(t-x)-\big{(}1+\varepsilon(x,y)\big{)}(t-y)\big{|}$
$=|c(c-1)(c-2)|\lambda_{x}^{\frac{1}{c-2}}\xi^{c-3}\big{|}(y-x)-\varepsilon(x,y)(t-y)\big{|}\text{.}$
Now we distinguish two cases. Firstly, we assume that $\varepsilon(x,y)\leq 2^{-\delta_{2}j-100}$. Note that
@@
and the same reasoning together with the fact that $\lambda_{y}\simeq 2^{r-cj}\simeq\lambda_{x}$ yields
@@
$2^{r-3j}\big{(}2^{(1-\delta_{2})j}-2^{-\delta_{2}j-100}2^{j-1}\big{)}\leq 2^{r-3j}\big{(}|y-x|-2^{-\delta_{2}j-100}|t-y|\big{)}\lesssim|f^{\prime\prime}(t)|\lesssim 2^{r-3j}2^{j}=2^{r-2j}\text{,}$
thus $2^{r-2j-\delta_{2}j}\lesssim|f^{\prime\prime}(t)|\lesssim 2^{r-2j}$. By van der Corput, see [11] page 208, for any interval $J\subseteq\{a_{x,y},\dotsc,b_{x,y}\}$ we get
$\bigg{|}\sum_{m\in J}e\left(\lambda_{x}|x-m|^{c}-\lambda_{y}|y-m|^{c}\right)\bigg{|}\lesssim 2^{\delta_{2}j}2^{\frac{r}{2}-j-\frac{\delta_{2}}{2}j}2^{j}+2^{-\frac{r}{2}+j+\frac{\delta_{2}}{2}j}=2^{\frac{\delta_{2}}{2}j}\big{(}2^{\frac{r}{2}}+2^{j-\frac{r}{2}}\big{)}\text{.}$
@@
$|K_{-,j}(x,y)|\lesssim 2^{-2j+\frac{\delta_{2}}{2}j+\frac{r}{2}}+2^{-j+\frac{\delta_{2}}{2}j-\frac{r}{2}}=2^{-j}\big{(}2^{-j+\frac{\delta_{2}}{2}j+\frac{r}{2}}+2^{\frac{\delta_{2}}{2}j-\frac{r}{2}}\big{)}$
and note that since $r\in\big{[}\nu j-2,(c+2\delta_{1})j+2\big{]}$, we get
$-j+\frac{\delta_{2}}{2}j+\frac{r}{2}\leq-j+\frac{\delta_{2}}{2}j+\frac{1}{2}(c+2\delta_{1})j+1=\frac{c-2+\delta_{2}+2\delta_{1}}{2}j+1$
@@
$\frac{\delta_{2}}{2}j-\frac{r}{2}\leq\frac{\delta_{2}}{2}j-\frac{\nu}{2}j+1\text{.}$
Remembering that $\delta_{2}<\min\big{\{}\frac{\nu}{100},\frac{2-c}{100}\big{\}}$ and $\delta_{1}<\frac{2-c}{4}$, we see that there exists $\rho^{\prime}>0$ such that $|K_{-,j}(x,y)|\lesssim 2^{-(1+\rho^{\prime})j}$.
Now let us assume that $\varepsilon(x,y)\geq 2^{-\delta_{2}j-100}$. Fix $\nu^{\prime}\in(\delta_{2},\min\{\frac{\nu}{2},\frac{2-c}{2}\})$ and define
$I_{x,y}\coloneqq\{m\in\{a_{x,y}\dotsc,b_{x,y}\}:\,|(y-x)-\varepsilon(x,y)(m-y)|\leq 2^{j-\nu^{\prime}j}\}\text{, and thus }|I_{x,y}|\lesssim 2^{(1+\delta_{2}-\nu^{\prime})j}\text{.}$
@@
$\Big{|}\sum_{m\in I_{x,y}}e\left(\lambda_{x}|x-m|^{c}-\lambda_{y}|y-m|^{c}\right)\psi^{-}_{j}(x-m)\psi_{j}^{-}(y-m)\Big{|}\lesssim 2^{-2j}|I_{x,y}|\lesssim 2^{-(1+\nu^{\prime}-\delta_{2})j}\text{,}$
and $11+\nu^{\prime}-\delta_{2}>$ 1 since $\nu^{\prime}>\delta_{2}$. We now treat $I^{1}_{x,y}$, and $I^{2}_{x,y}$ can be handled likewise. Let $c_{x,y}$ and $d_{x,y}$ be the left and right endpoints of $I^{1}_{x,y}$, and note that for any $t\in[c_{x,y},d_{x,y}]$ we have
$2^{r-3j}2^{j-\nu^{\prime}j}\lesssim|f^{\prime\prime}(t)|\lesssim 2^{r-3j}2^{j}\text{,}$
where for the first estimate we used the fact that if $t\in[c_{x,y},d_{x,y}]$, then $|(y-x)-\varepsilon(x,y)(t-y)|>2^{j-\nu^{\prime}j}$, and for the second estimate we used the fact that $\varepsilon(x,y)\lesssim 1$, since $\lambda_{x}\simeq\lambda_{y}$. We use van der Corput to conclude that for any subinterval $(I^{1}_{x,y})^{\prime}$ of $I^{1}_{x,y}$ we get
$\Big{|}\sum_{m\in(I^{1}_{x,y})^{\prime}}e\left(\lambda_{x}|x-m|^{c}-\lambda_{y}|y-m|^{c}\right)\Big{|}\lesssim 2^{\nu^{\prime}j}2^{\frac{r}{2}-j-\frac{\nu^{\prime}}{2}j}2^{j}+2^{-\frac{r}{2}+j+\frac{\nu^{\prime}}{2}j}=2^{\frac{\nu^{\prime}}{2}j+\frac{r}{2}}+2^{-\frac{r}{2}+j+\frac{\nu^{\prime}}{2}j}\text{,}$

### v2

@@ 5. Reducing the proof of Theorem 1 to the continuous case
and notice that by taking into account our choice of $M$, the estimate (2.7) and since $\varepsilon j=\frac{2-c}{4}j\geq 1$ and $\log(M)\leq\log(2^{\varepsilon j}+1)\leq\log(2^{\varepsilon j+1})\leq\log(2^{2\varepsilon j})\lesssim(2-c)j$, we get
(5.14) $\|L_{j}\|_{\ell^{1}}\lesssim 2^{-j}\sum_{2^{j-3}\leq n\leq 2^{j-1}}\min\bigg\{1,\frac{1}{M\|n^{c}\|}\bigg\}\lesssim(c-1)^{-\frac{1}{2}}2^{-j}j\varepsilon\big(2^{(1-\varepsilon)j}+2^{(c/2+\varepsilon/2)j}\big)\\ \lesssim(c-1)^{-\frac{1}{2}}\Big(j\varepsilon 2^{-\varepsilon j}+j\varepsilon 2^{\frac{c+\varepsilon-2}{2}j}\Big)\lesssim(c-1)^{-\frac{1}{2}}j\varepsilon 2^{-\varepsilon j}\lesssim(c-1)^{-\frac{1}{2}}2^{-\frac{\varepsilon}{2}j}\text{,}$
since $c+\varepsilon-2<-2\varepsilon$, because $\varepsilon<\frac{2-c}{3}$, and by (3.2). Thus $\|L_{j}\|_{\ell^{1}}\lesssim(c-1)^{-\frac{1}{2}}2^{-\frac{\varepsilon}{2}j}$. Now by defining $T_{j}f\coloneqq f*L_{j}$, we see that
$\|T_{j}\|_{\ell^{2}\to\ell^{2}}\lesssim(c-1)^{-\frac{1}{2}}2^{-\frac{\varepsilon}{2}j}\text{.}$
For $\tau\in\{+,-\}$, let
[…]
@@
(5.15) $\sum_{\mathsf{\tau}\in\left\{+,-\right\}}\sup_{\lambda\in X_{j}^{c}}\sum_{|m|\leq M}\frac{1}{|m|+1}\left|\sum_{n\neq 0}f(x-n)e((\lambda\mathsf{sign}^{\mathsf{i}}(n)+m)|n|^{c})\psi_{j}^{\tau}(n)\right|\\ \lesssim\sum_{\mathsf{\tau}\in\left\{+,-\right\}}\sup_{2^{(-c+\nu)j}<|\lambda|\leq M+1/2}\left|\sum_{n\neq 0}f(x-n)e(\lambda|n|^{c})\psi_{j}^{\tau}(n)\right|\log(M+1)\\ \lesssim(2-c)j\max_{\mathsf{\tau}\in\left\{+,-\right\}}\sup_{2^{(-c+\nu)j}<|\lambda|\leq 2^{2\varepsilon j}}\Big|\sum_{n\neq 0}f(x-n)e(\lambda|n|^{c})\psi_{j}^{\tau}(n)\Big|\text{,}$
since $\varepsilon j=\frac{2-c}{4}j\geq 1$ and $\log(M+1)\leq\log(2^{\varepsilon j}+1)\leq\log(2^{\varepsilon j+1})\leq\log(2^{2\varepsilon j})\lesssim(2-c)j$. Combining everything yields
@@
(5.16) $\Big\|\sup_{\lambda\in X^{c}_{j}}\big|\big(\hat{f}(\cdot)m_{j}^{\mathsf{i}}(\cdot,\lambda)\big)^{\vee}(x)\big|\Big\|_{\ell^{2}_{{\rm d}x}}\lesssim(2-c)j\max_{{\mathsf{\tau}\in\left\{+,-\right\}}}\bigg\|\sup_{2^{(-c+\nu)j}<|\lambda|\leq 2^{2\varepsilon j}}\Big|\sum_{n\neq 0}f(x-n)e(\lambda|n|^{c})\psi_{j}^{\tau}(n)\Big|\bigg\|_{\ell^{2}_{{\rm d}x}}\\ +(c-1)^{-\frac{1}{2}}2^{-\frac{\varepsilon}{2}j}\|f\|_{\ell^{2}}\text{,}$
@@
and we focus on the first term of the right hand side. We claim that it suffices to prove that for each $\tau\in\left\{+,-\right\}$ we have
(5.17) $\bigg\|\sup_{2^{\nu j}<|\lambda 2^{cj}|\leq 2^{(c+2\varepsilon)j}}\Big|\sum_{n\neq 0}f(x-n)e(\lambda|n|^{c})\psi_{j}^{\tau}(n)\Big|\bigg\|_{\ell^{2}_{{\rm d}x}}\lesssim(2-c)^{-1}j\big((2-c)(c-1)\big)^{-\frac{1}{4}}2^{-\frac{\varepsilon}{128}j}\|f\|_{\ell^{2}}\text{,}$
To see why this is the case note that once we establish (5.17) we may return to (5.16) to obtain
(5.18) $\Big\|\sup_{\lambda\in X^{c}_{j}}\big|\big(\hat{f}(\cdot)m_{j}^{\mathsf{i}}(\cdot,\lambda)\big)^{\vee}(x)\big|\Big\|_{\ell^{2}_{{\rm d}x}}\lesssim j^{2}\big((2-c)(c-1)\big)^{-\frac{1}{4}}2^{-\frac{\varepsilon}{128}j}\|f\|_{\ell^{2}}+(c-1)^{-\frac{1}{2}}2^{-\frac{\varepsilon}{2}j}\|f\|_{\ell^{2}}\\ \lesssim(2-c)^{-\frac{9}{4}}(c-1)^{-\frac{1}{4}}(\varepsilon j)^{2}2^{-\frac{\varepsilon}{128}j}\|f\|_{\ell^{2}}+(c-1)^{-\frac{1}{2}}2^{-\frac{\varepsilon}{2}j}\|f\|_{\ell^{2}}\lesssim(2-c)^{-\frac{9}{4}}(c-1)^{-\frac{1}{2}}2^{-\frac{\varepsilon}{256}j}\|f\|_{\ell^{2}}\text{,}$
where we took into account the fact that $\varepsilon=\frac{2-c}{4}$ and used the estimate $x^{2}2^{-x}\lesssim 2^{-x/2}$.
To establish (5.17) it will be important to have further control of the scale of $2^{cj}\lambda$, so we perform a further decomposition. Note that
(5.19) $\sup_{2^{\nu j}<\lambda 2^{cj}\leq 2^{(c+2\varepsilon)j}}\Big|\sum_{n\neq 0}f(x-n)e(\lambda|n|^{c})\psi_{j}^{\tau}(n)\Big|\\ \leq\sum_{r\in(2-c)\mathbb{Z}}\sup_{2^{\nu j}<\lambda 2^{cj}\leq 2^{(c+2\varepsilon)j}}\bm{1}_{[2^{r},2^{r+(2-c)})}(\lambda 2^{cj})\Big|\sum_{n\neq 0}f(x-n)e(\lambda|n|^{c})\psi_{j}^{\tau}(n)\Big|\\ \leq\sum_{\begin{subarray}{c}r\in(2-c)\mathbb{Z}\text{,}\\ \nu j-2\leq r\leq(c+2\varepsilon)j+2\end{subarray}}\sup_{2^{r}\leq\lambda 2^{cj}<2^{r+(2-c)}}\Big|\sum_{n\neq 0}f(x-n)e(\lambda|n|^{c})\psi_{j}^{\tau}(n)\Big|\text{,}$
where for the last equality we simply note that for any $r\in(2-c)\mathbb{Z}$ [^3] Here we use the standard notation $\lambda\mathbb{Z}\coloneqq\{\lambda k:\,k\in\mathbb{Z}\}$. Letting $r\in\mathbb{Z}$ instead of $r\in(2-c)\mathbb{Z}$ results in an exponential loss in the constants and since we try to optimize the constant the argument yields with respect to $c$, a finer control of the scale $2^{cj}\lambda$ is needed. such that $r<\nu j-2$ or $r>(c+2\varepsilon)j+2$ we have that
$\big[2^{r},2^{r+(2-c)}\big)\cap\big(2^{\nu j},2^{(c+2\varepsilon)j}\big]=\emptyset\text{.}$
@@
Note that the outer sum has $O((2-c)^{-1}j)$ terms and thus to establish (5.17) it suffices to prove that for any $r\in[\nu j-2,(c+2\varepsilon)j+2]$ we have that
(5.20) $\bigg\|\sup_{2^{r}\leq\lambda 2^{cj}<2^{r+1}}\Big|\sum_{n\neq 0}f(x-n)e(\lambda|n|^{c})\psi_{j}^{\tau}(n)\Big|\bigg\|_{\ell^{2}_{{\rm d}x}}\lesssim\big((2-c)(c-1)\big)^{-\frac{1}{4}}2^{-\frac{\varepsilon}{128}j}\|f\|_{\ell^{2}}\text{,}$
@@
We linearize the supremum; we will prove that for any measurable $\lambda_{x}\colon\mathbb{Z}\to[2^{r-cj},2^{r+(2-c)-cj})$ we have
(5.21) $\Big\|\sum_{n\neq 0}f(x-n)e(\lambda_{x}|n|^{c})\psi_{j}^{\tau}(n)\Big\|_{\ell^{2}_{{\rm d}x}}\lesssim\big((2-c)(c-1)\big)^{-\frac{1}{4}}2^{-\frac{\varepsilon}{128}j}\|f\|_{\ell^{2}}\text{.}$
We fix such a function and use the $TT^{*}$ method here. It is not difficult to see that if
@@
Fix $\rho\coloneqq\frac{\varepsilon}{64}=\frac{2-c}{256}$ and $\sigma\coloneqq\frac{\varepsilon}{16}=\frac{2-c}{64}$. There exists an absolute positive constant $C$ such that the following estimate holds
(5.22) $|K_{\tau,j}(x,y)|\lesssim\big((2-c)(c-1)\big)^{-\frac{1}{2}}2^{-(1+\rho)j}\bm{1}_{2^{(1-\sigma)j}\leq|x-y|\leq 2^{j}}+2^{-j}\bm{1}_{|x-y|\leq 2^{(1-\sigma)j}}\text{.}$
@@
We begin with certain straightforward reductions. Firstly, we assume that $y\geq x$ and $\tau=$- since symmetric arguments yield the result for the remaining cases. Secondly, we may assume that $|x-y|\geq 2^{(1-\sigma)j}$ since otherwise we clearly have
@@
Now for the case $y\geq x$ and $|x-y|\geq 2^{(1-\sigma)j}$ we note that
@@
since for the first sum we have that xmx-m and ymy-m have different signs and for the third one we note that xm>0x-m>0 and thus $\psi_{j}^{-}(x-m)=0$. We focus on bounding the last line of (5.23). For any fixed $x,y$ the phase of the exponential sum becomes
@@
where $a_{x,y},b_{x,y}$ are defined by $\{m\in\mathbb{Z}:\,m>y,\,2^{j-3}\leq|m-y|\leq 2^{j-1},\,2^{j-3}\leq|m-x|\leq 2^{j-1}\}\eqqcolon\{a_{x,y},\dotsc,b_{x,y}\}$. We see that
@@
(5.24) $|f^{\prime\prime}(t)|=|c(c-1)(c-2)|\xi^{c-3}\Big|\lambda_{x}^{\frac{1}{c-2}}(t-x)-\lambda_{y}^{\frac{1}{c-2}}(t-y)\Big|\\ =|c(c-1)(c-2)|\lambda_{x}^{\frac{1}{c-2}}\xi^{c-3}\bigg|(t-x)-\frac{\lambda_{y}^{\frac{1}{c-2}}}{\lambda_{x}^{\frac{1}{c-2}}}(t-y)\bigg|\text{,}$
@@
(5.25) $|f^{\prime\prime}(t)|=|c(c-1)(c-2)|\lambda_{x}^{\frac{1}{c-2}}\xi^{c-3}\big|(t-x)-\big(1+\varepsilon(x,y)\big)(t-y)\big|\\ =|c(c-1)(c-2)|\lambda_{x}^{\frac{1}{c-2}}\xi^{c-3}\big|(y-x)-\varepsilon(x,y)(t-y)\big|\text{.}$
Now we distinguish two cases. Firstly, we assume that $\varepsilon(x,y)\leq 2^{-\sigma j-100}$. Note that then
@@
and the same reasoning together with the fact that $\lambda_{y}^{\frac{1}{c-2}}\simeq\lambda_{x}^{\frac{1}{c-2}}$ yield
@@
$2^{r-3j}\big(2^{(1-\sigma)j}-2^{-\sigma j-100}2^{j-1}\big)\leq 2^{r-3j}\big(|y-x|-2^{-\sigma j-100}|t-y|\big)\lesssim\frac{|f^{\prime\prime}(t)|}{(c-1)(2-c)}\lesssim 2^{r-3j}2^{j}=2^{r-2j}\text{,}$
thus $2^{r-2j-\sigma j}\lesssim\frac{|f^{\prime\prime}(t)|}{(c-1)(2-c)}\lesssim 2^{r-2j}$. By van der Corput, see [11] page 208, for any interval $J\subseteq\{a_{x,y},\dotsc,b_{x,y}\}$ we get
(5.26) $\bigg|\sum_{m\in J}e\left(\lambda_{x}|x-m|^{c}-\lambda_{y}|y-m|^{c}\right)\bigg|\lesssim 2^{\sigma j}\big((c-1)(2-c)\big)^{\frac{1}{2}}2^{\frac{r}{2}-j-\frac{\sigma}{2}j}2^{j}+\big((c-1)(2-c)\big)^{-\frac{1}{2}}2^{-\frac{r}{2}+j+\frac{\sigma}{2}j}\\ \lesssim 2^{\frac{\sigma}{2}j}\Big(2^{\frac{r}{2}}+\big((c-1)(2-c)\big)^{-\frac{1}{2}}2^{j-\frac{r}{2}}\Big)\text{.}$
@@
$|K_{-,j}(x,y)|\lesssim 2^{-2j+\frac{\sigma}{2}j+\frac{r}{2}}+\big((c-1)(2-c)\big)^{-1/2}2^{-j+\frac{\sigma}{2}j-\frac{r}{2}}\lesssim\big((c-1)(2-c)\big)^{-1/2}2^{-j}\big(2^{-j+\frac{\sigma}{2}j+\frac{r}{2}}+2^{\frac{\sigma}{2}j-\frac{r}{2}}\big)\text{,}$
and note that since $r\in\big[\nu j-2,(c+2\varepsilon)j+2\big]$, we get
$j+\frac{\sigma}{2}j+\frac{r}{2}\leq-j+\frac{\sigma}{2}j+\frac{1}{2}(c+2\varepsilon)j+1=\frac{c-2+\sigma+2\varepsilon}{2}j+1$
@@
$\frac{\sigma}{2}j-\frac{r}{2}\leq\frac{\sigma}{2}j-\frac{\nu}{2}j+1\text{.}$
Remembering that $\sigma=\frac{\varepsilon}{16}$, $\nu=\frac{\varepsilon}{8}$ and $\varepsilon=\frac{2-c}{4}$, it is easy to check that the previous estimate yields that
(5.27) $|K_{-,j}(x,y)|\lesssim\big((c-1)(2-c)\big)^{-1/2}2^{-(1+\rho)j}\text{,}\quad\text{for $\rho=\frac{\varepsilon}{64}$\text{.}}$
It remains to treat the case when $\varepsilon(x,y)\geq 2^{-\sigma j-100}$. Fix $\nu^{\prime}\coloneqq\frac{3\sigma}{2}=\frac{3\varepsilon}{32}$ and define
$I_{x,y}\coloneqq\{m\in\{a_{x,y}\dotsc,b_{x,y}\}:\,|(y-x)-\varepsilon(x,y)(m-y)|\leq 2^{j-\nu^{\prime}j}\}\text{, and thus }|I_{x,y}|\lesssim 2^{(1+\sigma-\nu^{\prime})j}\text{.}$
@@
(5.28) $\Big|\sum_{m\in I_{x,y}}e\left(\lambda_{x}|x-m|^{c}-\lambda_{y}|y-m|^{c}\right)\psi^{-}_{j}(x-m)\psi_{j}^{-}(y-m)\Big|\lesssim 2^{-2j}|I_{x,y}|\lesssim 2^{-(1+\nu^{\prime}-\sigma)j}\lesssim 2^{-(1+\frac{\varepsilon}{64})j}\text{,}$
since $\nu^{\prime}-\sigma=\frac{\sigma}{2}=\frac{\varepsilon}{32}$.
We now treat $I^{1}_{x,y}$, and $I^{2}_{x,y}$ can be handled likewise. Let $c_{x,y}$ and $d_{x,y}$ be the left and right endpoints of $I^{1}_{x,y}$, and note that for any $t\in[c_{x,y},d_{x,y}]$ we have
$2^{r-3j}2^{j-\nu^{\prime}j}\lesssim\frac{|f^{\prime\prime}(t)|}{(c-1)(2-c)}\lesssim 2^{r-3j}2^{j}\text{,}$
where for the first estimate we used the fact that if $t\in[c_{x,y},d_{x,y}]$, then $|(y-x)-\varepsilon(x,y)(t-y)|>2^{j-\nu^{\prime}j}$, and for the second estimate we used the fact that $\varepsilon(x,y)\lesssim 1$, since
$\varepsilon(x,y)\coloneqq\Big(\frac{\lambda_{y}}{\lambda_{x}}\Big)^{\frac{1}{c-2}}-1\leq\Big(\frac{\lambda_{x}}{\lambda_{y}}\Big)^{\frac{1}{2-c}}\leq\bigg(\frac{2^{r+(2-c)-cj}}{2^{r-cj}}\bigg)^{\frac{1}{2-c}}\leq 2\text{.}$
We use van der Corput to conclude that for any subinterval $(I^{1}_{x,y})^{\prime}$ of $I^{1}_{x,y}$ we get
(5.29) $\Big|\sum_{m\in(I^{1}_{x,y})^{\prime}}e\left(\lambda_{x}|x-m|^{c}-\lambda_{y}|y-m|^{c}\right)\Big|\\ \lesssim\big((2-c)(c-1)\big)^{\frac{1}{2}}2^{\nu^{\prime}j}2^{\frac{r}{2}-j-\frac{\nu^{\prime}}{2}j}2^{j}+\big((2-c)(c-1)\big)^{-\frac{1}{2}}2^{-\frac{r}{2}+j+\frac{\nu^{\prime}}{2}j}\lesssim 2^{\frac{\nu^{\prime}}{2}j+\frac{r}{2}}+\big((2-c)(c-1)\big)^{-\frac{1}{2}}2^{-\frac{r}{2}+j+\frac{\nu^{\prime}}{2}j}\text{,}$